The potential at location A is 429 V. A positively charged particle is released there from rest and arrives at location B with a speed vB. The potential at location C is 783 V, and when released from rest from this spot, the particle arrives at B with twice the speed it previously had, or 2vB. Find the potential at B.
Answer #1


vA = 429 v

vC = 783 v

vB = ?


from A to B

using work energy theorem

work = change in KE

(vA-vB)*q = 0.5*m*vB^2 .......(1)

from C to B


at B speed = 2vB

(vC - vB) *q = 0.5*m*4vB^2........(2)

from 1 & 2


(vA-vB)*4 = vC - vB

4*429 - 4VB = 783 - vB

vb = 311 V    <<<<<------------answer